OpenStudy (anonymous):

Which ordered pair is a solution of the equation 4x - 2y = 16? a. (-5, -3) b. (-5, 3) c. (6, -4) d. (6, 4)

pooja195 (pooja195):

Plug each of your options in place of X and y to see what = s 16

pooja195 (pooja195):

\[\huge A.\] \[\huge4(-5) - 2(-3)= 16\]

pooja195 (pooja195):

\[\huge~B.\] \[\huge4x(-5)- 2(3) = 16\]

pooja195 (pooja195):

\[\huge~C.\] \[\huge 4(6)- 2(-4)= 16\]

pooja195 (pooja195):

\[\huge~D.\] \[\huge4(6) - 2(4) = 16\]

pooja195 (pooja195):

Im sure you can do the rest.